OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4 2 3 All real numbers No solution

OpenStudy (jhannybean):

Refer to the steps i've shown in your previous problem. http://openstudy.com/study#/updates/51c37ce1e4b069eb00c4c801

OpenStudy (anonymous):

it's No Soln\[-2( x+3)=-2(x-1)-4,---->-2x-6=-2x+2-4 and when u take -2x \to the \left \it would turn x \to 0\]